# On the handling of method failure in comparison studies

### v1

@@ 4.1.2 Design of the fictive study
In our fictive simulation study, we vary the underlying true OR (OR $\in\{2,3,4,5\}$) and exposure probability $p_{x}$ ($p_{x}\in\{0.25,0.5\})$ while keeping the sample size at the fixed value of $n.obs=50$. This results in eight distinct simulation scenarios (see Table 5 for an overview). For each scenario, we simulate 100 000100000100\,000100 000 outcome-exposure data sets (where the exposure variable $X$ is generated using a Bernoulli distribution with probability $P(X=1)=p_{x}$ and outcome $Y$ is generated using the logistic function with the linear effect of exposure $X$ specified as $\beta_{X}=\log(\text{OR})$). Five OR estimation methods are applied to the simulated data to obtain a point estimate for the true OR, namely “Fisher”, “Midp”, “Small” (all three implemented in R package epitools (Aragon et al., , 2020)), “Woolf” (Woolf et al., , 1955) (implemented in R package pairwiseCI (Schaarschmidt and Gerhard, , 2019)), and “Manual”, which corresponds to the manual computation $\widehat{OR}=\frac{n_{11}\cdot n_{00}}{n_{01}\cdot n_{10}}$. Note that we omit method “Wald” from the epitools package since its estimation output is identical to the manual computation (i.e., method “Manual”). From the five estimators under investigation, only Small and Woolf can handle sampling zeros internally. For each OR estimator and simulation scenario, bias on the $\log$-scale is computed as
[…]
and the MSE on the $\log$-scale is obtained as
[…]
Note that the $\log$-transformation ensures symmetry around the value of “no association”. For both performance measures, values closer to 00 indicate better performance (Morris et al., , 2019).
4.1.3 Occurrence and handling of method failure
@@
Our analysis consists of two parts. First, we compare the performance of the OR estimation methods regarding performance measures bias and MSE when ignoring data sets with sampling zeros for all methods versus ignoring data sets only for the failing methods (“ad hoc approaches”). While we are aware that users often apply the Haldane-Anscombe correction when encountering sampling zeros, this first part illustrates how the preference of some methods may mainly be driven by the way data sets with sampling zeros are discarded.
Second, we repeat the analysis implementing the principle of fallback strategies suggested in Section 3.3.5. As fallbacks for methods Midp and Fisher for data sets with sampling zeros, we (i) either apply the Haldane-Anscombe correction, (ii) use method Small, or (iii) use method Woolf. For the manual estimation of the OR (method “Manual”), we only apply the Haldane-Anscombe correction as a fallback, assuming that a user applying manual computation in the first place, which is very easy to implement, is unlikely to resort to a method that requires notably more effort. Together with Woolf and Small, nine “methods” (“pure” methods or pipelines) are compared in total and the results are presented in Section 4.1.4.
Note that since we focus on the effect of handling method failure (instead of the actual results of the comparison), and this effect becomes clearer through examining the individual scenarios, we do not aggregate the performances across the eight scenarios in either of the two analysis parts.
@@
Figure 1: Performance ranks of the OR estimation methods across the eight simulation scenarios, derived from the empirical bias in subfigure (A) and from the MSE in subfigure (B). In both subfigures, the ranks are compared between discarding data sets yielding failure for the failing methods only (denoted by “Failing” on the $x$-axis) versus for all methods (denoted by “All” on the $x$-axis). A rank of 1 means that the method has the lowest empirical bias (A) or MSE (B) in the given scenario. Note that Small and Woolf are the only methods in this selection that can deal with sampling zeros internally. Abbreviation of methods: “Fis”: Fisher; “Mid”: Midp“; “Sma”: Small; “Woo”: Woolf; “Man”: Manual (i.e., manual estimation of the OR).
@@
Ad hoc approaches  For the first analysis, Figure 1 shows that the way of discarding data sets with sampling zeros can substantially influence the results of the method comparison. This particularly affects estimator Woolf, one of the two methods that can handle sampling zeros internally. In terms of bias (Figure 1 (A)), its performance rank deteriorates in most scenarios and by up to three positions when discarding data sets with sampling zeros for all methods (including Woolf and Small). The reason for this is that across the simulated data sets without sampling zeros, Woolf slightly underestimates the true OR, while for the few data sets with sampling zeros, Woolf notably overestimates it. This leads to top performance when including all data sets in its performance assessment (while discarding data sets with sampling zeros for the failing methods). Even more so, the manner of discarding data sets with sampling zeros affects Woolf’s performance rank regarding MSE (Figure 1 (B)). In some scenarios, Woolf’s rank improves from bottom to top rank, indicating its poorer performance for data sets with sampling zeros. The performance rank of Small, the second method that can handle sampling zeros internally, is less affected by this for both bias and MSE. In particular, it scores last in terms of bias across all scenarios and for both manners of discarding (Figure 1 (A)). However, its performance rank changes by up to two positions for MSE (Figure 1 (B)). The performance ranks of the remaining methods are also affected regarding bias and MSE, where the manner of discarding data sets with sampling zeros affects their performance ranks by up to two positions.
@@
Note that the manner of discarding data sets has a larger impact on the performance comparison for scenarios with a higher proportion of sampling zeros. In particular, scenarios 5 and 6, with a maximum of $0.11\%$ of data sets containing sampling zeros, yield equal performance rankings across both handlings. However, differences in rankings within a given scenario are already observable with a proportion of $0.5\%$ of data sets with sampling zeros.
Figure 2: Performance ranks of the OR estimation methods across the eight simulation scenarios when applying fallback strategies to methods Midp (“Mid”), Fisher (“Fis”), and Manual (“Man”). The rankings are derived from the empirical bias in subfigure (A) and from the MSE in subfigure (B). A rank of 1 means that the method has the lowest empirical bias (A) or MSE (B) in the given scenario. Note that for Midp and Small, the fallbacks Small (“Sma”), Woolf (“Woo”), and the Haldane-Anscombe correction (“$+0.5$”) are used. For Manual, only the Haldane-Anscombe correction is used as fallback. Small and Woolf are the only methods in this selection that can deal with sampling zeros internally, therefore not requiring any fallback. This amounts to nine “methods” in total. The combination of original method $m_{1}$ and fallback $m_{2}$ is marked by “$m_{1}/m_{2}$”.
4.1.4 Recommended handling
@@
As expected, the relative performance between Woolf and Small remains the same between applying fallbacks for the other methods (Figure 2) versus discarding data sets with sampling zeros for the failing methods only (“failing” parts in Figure 1 (A) and (B)). This makes sense since neither method requires a fallback, resulting in unchanged bias and MSE. Midp’s overall good performance regarding bias and MSE is underpinned by using fallbacks Small and the Haldane-Anscombe correction. Here, it is debatable which fallback is preferable. Both fallbacks for Midp frequently switch ranking positions across the performance measures and simulation scenarios. However, fallback Small might be even easier to implement than the Haldane-Anscombe correction, requiring only a single parameter modification from Midp. Note the interesting observation that Small scores below the remaining methods regarding bias when applied alone, but performs well as a fallback for Midp.
@@ 4.2.1 General
The fictive study aims to compare the performance of two methods for constructing confidence intervals (CIs) for the generalization performance of a prediction model based on resampling-based estimates, such as cross-validation. The “naive” method “$N$” for constructing the CIs ignores the dependence between the generalization performances in the resampling repetitions, while method “$C$” corrects for it. The performance of both methods is assessed in terms of coverage, which is the probability that a constructed CI includes the true value of the generalization performance. It should be close to its nominal value, e.g. $0.95$ if we consider $95\%$-CIs.
@@ 4.2.2 Design of the fictive study
Classification trees (Breiman et al., , 1984) are considered as models for predictive modeling. As benchmark data set, we use the data set provided by the SAPA project (Condon et al., , 2017), further processed according to Klau et al., (2023). The response variable is binary and five variables from the data set are used as predictors. The generalization performance is quantified using the Area Under the Curve (AUC), where a value of $0.5$ corresponds to a random prediction and a value of $1$ indicates perfect classification. The AUC is estimated using $15$ repetitions of repeated subsampling ($i=1,\ldots,15$) with a 4:1 ratio (i.e., the data is randomly split $15$ times into $80\%$ on which the model is trained and $20\%$ on which the AUC is obtained (Hastie et al., , 2009)).
@@
where $\overline{\widehat{AUC}}$ represents the mean AUC estimate over the $15$ subsampling repetitions and $S^{2}_{\widehat{AUC}_{i}}$ the corresponding sample variance, while $t_{0.975,14}$ is the $95\%$-quantile of the $t$-distribution with $15-1=14$ degrees of freedom. The fixed correction term $c>0$ incorporates the correlation structure between the subsampling repetitions (Nadeau and Bengio, , 2003). Method $N$, on the other hand, does not apply any correction, i.e. sets $c=0$ in the above formula.
[…]
The study is conducted using R package mlr3, method $N$ is implemented using function t.test() from package base and for method $C$, $\overline{\widehat{AUC}}$ and $S^{2}_{\widehat{AUC}_{i}}$ are obtained using functions mean() and sd() (from the same package).
@@
4.2.3 Occurrence and handling of method failure
@@
    for method $N$, set “undefined” confidence intervals to “not covering the true AUC”.
@@
The coverages resulting from the ad hoc approaches are displayed in Table 6 (a). While the coverage of method $C$ is constant across all three ad hoc approaches (coverages above the nominal level indicate conservativeness), method $N$, while too liberal across all approaches, is even more liberal using the ad hoc imputation approach.
@@ 4.2.4 Recommended handling
We now closely inspect the modeling processes underlying the study. The goal is to find out the root cause of the error message and to adapt the handling of the undefined performance values accordingly. The errors in method $N$ occur when the estimated AUCs are equal across all $15$ subsampling repetitions, resulting in a sample variance $S^{2}_{\widehat{AUC}{i}}$ of 0. This zero variance in the denominator of the $t$-statistic causes an error in function t.test(). However, focusing on the confidence interval, we realize that from a theoretical point of view, a sample variance of 0 simply means that the CI is the zero-width interval $\left\{\overline{\widehat{AUC}}\right\}$. Therefore, method $N$ not yielding a CI (and returning an error) is a technical artifact of its implementation. A straightforward solution would be to modify the implementation of method $N$, using mean() and sd() instead of t.test(), so that it computes the zero-width CI, which can enter the calculation of coverage as any other CI (see Table 6 (b) for the resulting coverage).
However, taking the investigation another step further reveals that a sample variance $S^{2}_{\widehat{AUC}{i}}=0$ always occurs alongside all $15$ estimated (and the approximate true) AUC being equal to $0.5$. Ultimately, this worst possible performance of the classification tree (“random prediction”) occurs when the classification tree does not perform any splitting. This is related to the characteristics of the data set. For example, the problem disappears if we include additional predictors: splitting then occurs, leading to an AUC higher than $0.5$.
The adequate handling of method failure in this illustrative study, as described in Section 3, depends on the aim of the study. In particular, a prediction problem in which splitting often does not take place can be viewed as an extreme scenario. If the study’s scope includes such extreme scenarios, one could attempt to suitably modify parameters of the classification tree to favor splitting. If, however, the study aims to compare both methods in regular (i.e., less extreme) scenarios, one might consider additional predictors to favor splitting more strongly, resort to an alternative prediction model (e.g., logistic regression), an alternative measure of generalization performance (e.g., accuracy), or use a different benchmarking data set.
Finally, note that this example illustrates that method failure may be complex and that the interplay causing it may not be immediately identifiable, even if the benchmark study consists of a single scenario only. In a real benchmark study which typically contains multiple scenarios, thorough investigations can become even more difficult.
@@ 5 Discussion
Method failure in comparison studies, often manifesting as issues regarding calculation (e.g., non-convergence), runtime, or memory issues, complicates the desired comparison of unconditional performance between the methods under investigation, i.e. across all simulated or benchmark data sets. Our informal literature review and illustrations both make clear that authors should not blindly rely on the initially perceived manifestation of failure but rather make extensive investigations of the sources of failure, which often guides the choice of the handling of the resulting undefined values.
@@
Even if we argue for decisions on a case-by-case basis after careful inspection of the failures, we show and illustrate that some general principles hold in most situations. In particular, method failure usually leads to non-existing values instead of just “missing”, i.e., existent but unobserved values. This means imputation, a common but missing-data-inspired approach to handling method failure, is often inadequate. Another popular and seemingly logical approach, namely discarding data associated with failure (either for all or failing methods), almost never enables the comparison of unconditional performance. In this work, we provide alternative approaches that might be more appropriate to most occurrences of method failure.
@@
An approach that in our view deserves more attention than it receives currently in the literature consists in evaluating whole “pipelines” including one or several fallback strategies in case of method failure. This approach reflects the typical behavior of most users of methods in practice while enabling the comparison of unconditional performance if constructed carefully: if one method does not work, they resort to an alternative one. Future work is required to collect experiences and define standards regarding such pipelines and their evaluation. Above all, transparent reporting of the frequency of failure and of the corresponding handling is always recommended.
@@
A limitation of our recommendations is that, beyond obvious cases, they typically require both high time resources and thorough expertise in the underlying modelling processes and methods. Researchers conducting comparison studies may not always have these resources to the full extent. This particularly applies to a nuanced examination of the interactions between methods and data sets leading to method failure in different instances. We understand that this bottleneck may impair the full implementation of our recommendations in practice, hence our attempt to provide a slimmed-down version of the recommendations. When time resources and expertise are restricted, transparent reporting is all the more important. Also, our viewpoint on method failure as leading to “undefined” performance values does not always hold unrestrictedly, which can complicate the implementation of our recommendations. However, with method failure being a complex matter in and of itself, this underscores the importance of a case-by-case evaluation, as described above.
@@
In this article, method failure has been treated as black-or-white. However, in practice, whether or not a method has failed is itself a value of judgment. One such example is near-separation (van Smeden et al., , 2016), which does not necessarily make itself obvious, hence the existence of dedicated methods to detect it (Mansournia et al., , 2017). Such grey-area failures add further complexities to benchmark and simulation studies: overall performance may depend on the specific criteria used to judge failure.
@@
To sum up, we believe there is not “one” right way to handle method failure in all situations. Instead, it is important to carefully choose an approach depending on the context, and then report, explain (and ideally discuss) the decisions made. With our work, we hope to contribute meaningful ideas on how to think about method failure. They lead to helpful strategies for dealing with it, as outlined as part of our recommendations. However, we also believe that undefined performance values are only one of many issues related to comparison studies that deserve much more time, efforts and methodological strength towards more reliable evidence on the behavior of methods in the context of a “replication crisis in methodological research” (Boulesteix et al., , 2020). We need more well-designed neutral comparison studies (Boulesteix et al., , 2017)—late phase studies according to the terminology of Heinze et al., (2024). In such studies, no time is spent on the development of new methods. This leaves time for implementing our recommendations and diving into the technical details of the methods under investigation.

### v2

@@ 4.1.2 Design of the fictive study
We vary the underlying true OR (OR $\in\{2,3,4,5\}$) and exposure probability $p_{x}$ ($p_{x}\in\{0.25,0.5\})$ while keeping the sample size fixed at $n.obs=50$. This results in eight distinct simulation scenarios (see Table 4 for an overview). For each scenario, we simulate 100 000100000100\,000100 000 outcome-exposure data sets. Five OR estimation methods are applied to the simulated data to obtain a point estimate for the true OR. “Manual” corresponds to the manual and straightforward computation $\widehat{OR}=\frac{n_{11}\cdot n_{00}}{n_{01}\cdot n_{10}}$. The remaining four methods, namely “Fisher,” “Midp,” “Small” (all three implemented in R package epitools (Aragon et al., , 2020)), and “Woolf” (Woolf et al., , 1955) (implemented in R package pairwiseCI (Schaarschmidt and Gerhard, , 2019)) employ more complex modeling strategies. From the five estimators under investigation, only Small and Woolf can handle sampling zeros internally. For each simulation scenario, the estimators are compared regarding bias on the $\log$-scale ($\log$-transformation ensuring symmetry around the value of “no association”).
Our analysis consists of two parts. First, we compare the performance of the OR estimation methods when discarding data sets with sampling zeros for all methods versus only for the failing methods (“ad hoc approaches”; Section 4.1.3). Though users often apply the Haldane-Anscombe correction when encountering sampling zeros, this first part illustrates how the preference of some methods may mainly be driven by the way data sets with sampling zeros are discarded.
Second, we repeat the analysis implementing the principle of fallback strategies suggested in Section 3.1.5. This includes important considerations that must be made when choosing suitable fallbacks. For instance, for the manual estimation of the OR (method “Manual”), only the Haldane-Anscombe correction is applied as a fallback, assuming that a user applying the manual computation in the first place, which is very easy to implement, is unlikely to resort to a method that requires notably more effort. On the other hand, for methods Midp and Fisher for data sets with sampling zeros, we (i) either apply the Haldane-Anscombe correction, (ii) use method Small, or (iii) use method Woolf, assuming that corresponding users are more likely to employ fallbacks that are slightly more difficult to apply. Together with Woolf and Small, this leads to nine “methods” (“pure” methods or pipelines) being compared in total. The results are presented in Section 4.1.4.
4.1.3 Occurrence of method failure and ad hoc handlings
@@
Figure 1: Performance ranks of the OR estimation methods across the eight simulation scenarios based on empirical bias on the $\log$-scale, where a rank of 1 means that the method has the lowest empirical bias in the given scenario. (A) Ranks are compared between discarding data sets for the individual failing methods only (“Single” on the $x$-axis) versus for all methods (“All” on the $x$-axis). (B) The ranks are compared after applying fallback strategies to methods Midp, Fisher, and Manual. For Midp and Small, the fallbacks Small, Woolf, and the Haldane-Anscombe correction (“$+0.5$”) are used. For Manual, only the Haldane-Anscombe correction is used as fallback. Small and Woolf can deal with sampling zeros internally.
Abbreviation of methods: “Fis”: Fisher; “Mid”: Midp“; “Sma”: Small; “Woo”: Woolf; “Man”: Manual (i.e., manual estimation of the OR). (B): The combination of original method $m_{1}$ and fallback $m_{2}$ is marked by “$m_{1}/m_{2}$.”
@@
Ad hoc approaches  For the first analysis, Figure 1(A) shows that the way of discarding data sets with sampling zeros can substantially influence the results of the method comparison. This can lead to changes in the performance rank by up to three positions in some scenarios (see method Woolf) and particularly affect which method is chosen as the best. Note that a more detailed investigation of Woolf’s performance shows that for the data sets without sampling zeros, Woolf slightly underestimates the true OR, while for the few data sets with sampling zeros, Woolf notably overestimates it. This leads to top performance when all data sets are included in its performance assessment (while discarding data sets with sampling zeros for the failing methods).
@@
Generally, the manner of discarding data sets has a larger impact on the performance comparison for scenarios with a higher proportion of sampling zeros. However, differences in rankings within a given scenario are already observable with a proportion of $0.5\%$ of data sets with sampling zeros.
4.1.4 Recommended handling: using fallbacks
@@
Woolf’s overall good performance across all scenarios is underpinned when using fallbacks for all methods affected by method failure. Furthermore, the comparison between Woolf and Midp, another method performing well in most scenarios, becomes clearer through the use of fallbacks. While they often alternate in rank (and especially between rank 1 and 2) when removing data sets (Figure 1 (A)), their performance comparison is constant across all scenarios when a fallback is applied to Midp (Figure 1 (B)). Since this applies to both fallbacks, the Haldane-Anscombe correction and Small, it is debatable which fallback is preferable. Both fallbacks for Midp frequently switch ranking positions across the performance measures and simulation scenarios, however, fallback Small might be even easier to implement than the Haldane-Anscombe correction, requiring only a single parameter modification from Midp.
@@ 4.2.1 General
The fictive study aims to compare the performance of two methods for constructing confidence intervals (CIs) for the generalization performance of a prediction model based on resampling-based estimates, such as cross-validation. The “naive” method “$N$” for constructing the CIs ignores the dependence between the generalization performances in the resampling repetitions, while method “$C$” corrects for it. The performance of both methods is assessed in terms of coverage. Note that a comprehensive benchmark study of methods for constructing CIs for the generalization error is performed by Schulz-Kümpel et al., (2024).
@@ 4.2.2 Design of the fictive study
Classification trees (Breiman et al., , 1984) are considered as models for predictive modeling and the data set provided by the SAPA project (Condon et al., , 2017), further processed according to Klau et al., (2023), is used as the benchmark data set. The response variable is binary. The generalization performance is quantified using the Area Under the Curve (AUC). The AUC is estimated using $15$ repetitions of repeated subsampling ($i=1,\ldots,15$) with a 4:1 ratio (i.e., the data is randomly split $15$ times into $80\%$ on which the model is trained and $20\%$ on which the AUC is obtained (Hastie et al., , 2009)).
@@
where $\overline{\widehat{AUC}}$ represents the mean AUC estimate and $S^{2}_{\widehat{AUC}_{i}}$ the sample variance over the $15$ subsampling repetitions. $t_{0.975,14}$ is the corresponding quantile of the $t$-distribution. A fixed correction term $c>0$ incorporates the correlation structure between the subsampling repetitions for method $C$ (see Nadeau and Bengio, (2003) for the construction of c). Method $N$, on the other hand, sets $c=0$ in the above formula.
To estimate the coverages of methods $N$ and $C$, the following procedure is repeated for $1000$ iterations (see Figure 2). The SAPA data set is randomly split into a large test set $D_{test}$ ($80\%$ of observations) and a training set $D_{train}$ (20% of observations). The true AUC is approximated by applying the trained model to $D_{test}$. The corresponding estimated AUC, on the other hand, is obtained via repeated subsampling on $D_{train}$. Confidence intervals for the true AUC are constructed according to Equation  (1) with corresponding values of $c$ for methods $N$ and $C$. Finally, the estimated coverages of methods $N$ and $C$ are the proportions of the $1000$ iterations where the respective confidence intervals cover the approximated true AUC.
The study is conducted in R using package mlr3, method $N$ is implemented using function t.test() from package base and for method $C$, $\overline{\widehat{AUC}}$ and $S^{2}_{\widehat{AUC}_{i}}$ are obtained using functions mean() and sd() (from the same package).
@@
4.2.3 Occurrence of method failure and ad hoc handlings
@@
    for method $N$, set “undefined” confidence intervals to “not covering the true AUC.”
@@
The coverages resulting from the ad hoc approaches are displayed in Table 5 (a). While the coverage of method $C$ is constant across all three ad hoc approaches, method $N$, while too liberal across all approaches, is even more liberal using the ad hoc imputation approach.
@@ 4.2.4 Recommended handling
We now closely inspect the modeling processes underlying the study to find the root cause of the error message and to adapt the handling of method failure accordingly.  Investigations show that the errors in method $N$ occur when the estimated AUCs are equal across all $15$ subsampling repetitions, resulting in a sample variance $S^{2}_{\widehat{AUC}{i}}$ of 0. This zero variance in the denominator of the $t$-statistic causes an error in function t.test(). Focusing on the confidence interval, however, a sample variance of 0 simply implies a zero-width CI $\left\{\overline{\widehat{AUC}}\right\}$. Therefore, method $N$ not yielding a CI (and returning an error) is a technical artifact of its implementation. A straightforward solution would be to modify the implementation of method $N$, using mean() and sd() instead of t.test(), so that it computes the zero-width CI, which can enter the calculation of coverage as any other CI (see Table 5 (b) for the resulting coverage).
However, upon yet further investigations, it becomes clear that a sample variance $S^{2}_{\widehat{AUC}{i}}=0$ always coincides with all 15 estimated (and the approximate true) AUC being $0.5$. Ultimately, this worst possible performance of the classification tree (“random prediction”) occurs when the classification tree does not perform any splitting. This is related to the characteristics of the data set. For example, the problem disappears if we include additional predictors: splitting then occurs, leading to an AUC higher than $0.5$.
This illustrates well that the further the investigations are taken, the apparent adequate handling of method failure can change notably. In this case, based on the information gained, the adequate handling depends on the aim of the study and requires further considerations such as the following. A prediction problem where splitting often does not occur can be viewed as an edge case. If the study’s scope includes such edge cases, one could attempt to suitably modify the parameters of the classification tree to favor splitting. If, however, the study aims to compare both methods in regular scenarios, one might consider adding additional predictors to favor splitting more strongly, resort to an alternative prediction model (e.g., logistic regression), an alternative measure of generalization performance (e.g., accuracy), or use a different benchmarking data set.
@@ 5 Discussion
Method failure in comparison studies, often manifesting as issues regarding calculation (e.g., non-convergence), runtime, or memory issues, complicates the desired comparison of unconditional performance between the methods under investigation, i.e., across all simulated or benchmark data sets. Our informal literature review and illustrations both make clear that authors should not blindly rely on the initially perceived manifestation of failure but rather make extensive investigations of the underlying factors causing the failure, which often guides the choice of the handling of the resulting undefined values.
@@
Even if we argue for decisions on a case-by-case basis after careful inspection of the failures, we show and illustrate that some general principles hold in most situations. In particular, method failure usually leads to non-existing values instead of just “missing,” i.e., existent but unobserved values. This means imputation, a common but missing-data-inspired approach to handling method failure, is usually inadequate. The same applies to the popular and seemingly logical approach of discarding data associated with failure (either for all or failing methods), which almost never enables the comparison of unconditional performance. In this work, we provide alternative approaches that might be more appropriate to most occurrences of method failure.
@@
An approach that in our view deserves more attention than it receives currently in the literature consists of evaluating whole “pipelines,” including one or several fallback strategies in case of method failure. This approach reflects the typical behavior of most users of methods in practice while enabling the comparison of unconditional performance if constructed carefully: if one method does not work, they resort to an alternative one. Future work is required to collect experiences and define standards regarding such pipelines and their evaluation. Above all, transparent reporting of the frequency of failure and of the corresponding handling is always recommended.
@@
A limitation of our recommendations is that, beyond obvious cases, they typically require both high time resources and thorough expertise in the underlying modelling processes and methods. Researchers conducting comparison studies may not always have these resources to the full extent. This particularly applies to a nuanced examination of the interactions between methods and data sets leading to method failure in different instances. We understand that this bottleneck may impair the full implementation of our recommendations in practice, hence our attempt to provide a slimmed-down version of the recommendations. When time resources and expertise are restricted, transparent reporting is all the more important. Furthermore, this underlines the necessity for openly sharing code, data, and ideally intermediate results, as this allows other researchers to explore the observed instances of method failure in greater detail.
Also, our viewpoint on method failure as leading to “undefined” performance values does not always hold unrestrictedly, which can complicate the implementation of our recommendations. However, with method failure being a complex matter in and of itself, this underscores the importance of a case-by-case evaluation, as described above.
@@
In this article, method failure has been treated as black-or-white. However, in practice, whether or not a method has failed is itself a value of judgment. One such example is near-separation (van Smeden et al., , 2016), which does not necessarily make itself obvious, hence the existence of dedicated methods to detect it (Mansournia et al., , 2017). Such gray-area failures add further complexities to benchmark and simulation studies: overall performance may depend on the specific criteria used to judge failure.
Partial failure of a method is also possible. For example, in a simulation study, a method might return a valid point estimate but not its standard error, or a point estimate for one estimand but not another. Many of the principles discussed in this article apply in such situations, for example, the principle to avoid imputing estimates where a method has failed, and the notion of a pipeline (when the standard error estimator fails, revert to a backup estimator if available). There are, however, further subtleties. For example, it may be inadvisable for the fallback approach to keep the point estimate from method A and fallback on the standard error estimated from method B.
@@
To sum up, we believe there is not “one” right way to handle method failure in all situations. Instead, it is important to carefully choose an approach depending on the context, and then report, explain (and ideally discuss) the decisions made. With our work, we hope to contribute meaningful ideas on how to think about method failure. They lead to helpful strategies for dealing with it, as outlined as part of our recommendations. However, we also believe that method failure is only one of many issues related to comparison studies that deserve much more time, efforts, and methodological strength towards more reliable evidence on the behavior of methods in the context of a “replication crisis in methodological research” (Boulesteix et al., , 2020). We need more well-designed neutral comparison studies (Boulesteix et al., , 2017)—late phase studies according to the terminology of Heinze et al., (2024). In such studies, no time is spent on the development of new methods. This leaves time for implementing our recommendations and diving into the technical details of the methods under investigation.